### v1

@@ Abstract
The strong magnetic fields of neutron stars are closely linked to their observed thermal, spectral, and timing properties, such as the distribution of spin periods and their derivatives. To understand the evolution of astrophysical observables over time, it is essential to develop robust theoretical frameworks and numerical models that solve the coupled thermal and magnetic field evolution equations, incorporating detailed microphysical inputs like thermal and electrical conductivities and neutrino emission rates. These efforts are key to uncovering how the strength and geometry of magnetic fields change with age, ultimately shedding light on the diverse phenomenology of neutron stars. In this review, we outline the fundamental theory underlying magneto-thermal evolution models, with an emphasis on numerical methods and a comprehensive set of benchmark tests intended to guide current and future code development. We revisit established results from axisymmetric simulations, highlight recent progress in fully three-dimensional models, and offer a perspective on the anticipated developments in this rapidly evolving field.
Keywords:
Neutron stars Pulsars Late stages of stellar evolution Magnetic fields Numerical simulations
@@ 1 Introduction
Regarding observation, NSs were first discovered as rotation-powered radio pulsars (standing for pulsating stars, due to their periodic signal). These so-called standard pulsars constitute the most numerous class of known NSs, approaching four thousand identified members. [^1] See the online Australia Telescope National Facility pulsar catalog, http://www.atnf.csiro.au/research/pulsar/psrcat/ The number is continuously increasing thanks to the progress in wide-field low-frequency radio surveys and the use of the new generation of high-sensitivity radio interferometers, like LOFAR (van Haarlem, 2013) and the soon-available Square Kilometre Array (SKA). The latter is foreseen to be able to detect many thousands of regular pulsars.
@@
To a lesser extent, NSs have also been observed in $X$-rays (about one hundred NSs so far), as persistent or transient sources, and/or as $\gamma$-ray pulsars (340 in the recent third Fermi-LAT pulsar catalog, Smith et al 2023). The origin of this high-energy radiation is typically non-thermal, originated by particle acceleration (synchro-curvature emission, Zhang and Cheng 1997; Viganò et al 2015) or Compton up-scattering of lower-energy photons by the particles composing the magnetospheric plasma (Lyutikov and Gavriil, 2006). An exception is the soft X-ray thermal emission from the surface, observed in only a few dozen, mostly young, neutron stars Potekhin et al (2015a); despite its rarity, it is highly relevant to this review.
@@
A particularly intriguing class of isolated NSs are the magnetars (Mereghetti et al, 2015; Turolla et al, 2015; Kaspi and Beloborodov, 2017; Rea and De Grandis, 2025), relatively slow rotators with typical spin periods of several seconds and ultra-strong timing-inferred magnetic fields ($10^{13}$$10^{15}$ G). In most cases, they show a relatively high persistent (i.e., constant over many years) X-ray luminosity ($L_{x}\approx 10^{33}$$10^{35}$ erg/s), well exceeding their rotational energy losses, in contrast with radio (standard) and $\gamma$-ray pulsars. This leads to the conclusion that the main source of energy is provided by the strong magnetic field, instead of rotational energy. Magnetars are also identified for their complex transient phenomenology in high energy $X$-rays and $\gamma$-rays, including short (tenths of a second) bursts, occasional energetic outbursts with months-long afterglows (Rea and Esposito, 2011; Coti Zelati et al, 2018) and, much more rarely (only three observed so far), giant flares (Hurley et al, 1999; Palmer et al, 2005). During giant flares, the energy release is as large as $10^{46}$ erg in less than a second. The source of energy of such transient, violent behavior is also generally agreed to be of magnetic origin, as originally proposed in Thompson and Duncan 1995, 1996.
@@
Although isolated NSs have been historically differentiated in sub-classes, mostly based on observational grounds (detectability in $X$ and/or radio, transient vs. persistent properties, and presence/absence of pulsations), there is arguably no sharp boundary between classes, and the distributions of their physical properties, such as the inferred magnetic field, partially overlap. Indeed, the evidence accumulated in the last few decades has shown that the presence of a strong dipolar field, which can be reliably estimated from timing properties (see Sect. 5), is not a sufficient condition to trigger observable magnetar-like events. In contrast, a growing number of NSs with relatively low inferred surface dipolar magnetic fields have been observed showing magnetar-like activity. These include some normal pulsars (Rea et al, 2010, 2012, 2013, 2014; Gavriil et al, 2008; Göğüs et al, 2016; Lower et al, 2021; Uzuner et al, 2023), often referred to as low-field magnetars, but also other sources belonging to the sub-class of central compact objects (CCOs), a handful of young NSs surrounded by a supernova remnant, detectable due to a persistent, mostly non-pulsating X-ray emission (De Luca, 2017). It is now evident that the complex, non-linear dynamics of the internal magnetic field, coupled with its interaction with the magnetospheric plasma, is crucial for understanding the observed phenomena.
@@ 2 Neutron star cooling
The evolution of the temperature in a NS was theoretically explored even before the first detections, in the 1960s (Tsuruta, 1964). Today, NS cooling is the most widely accepted terminology for the research area studying how internal and surface temperature evolve as NSs age and their observable effects. We refer the interested reader to the introduction in the review by Potekhin et al (2015b) for a thorough historical overview of the foundations of the NS cooling theory.
@@
The high thermal conductivity of the core rapidly leads to a nearly isothermal [^3] We employ the term ”isothermal” in the relativistic sense, including metric corrections, as we describe below. state within the first year. After a few years, the gradual long-term cooling of residual heat proceeds slowly, on a timescale of $10^{5}-10^{6}$ years, with temperature gradients essentially limited to the crust and envelope. These temperature gradients are primarily radial, but, as discussed in Sect. 2.2, they can also develop in the angular directions in the presence of strong magnetic fields, leading to surface temperature variations that influence the observed X-ray spectra. Notably, the absence of strong temperature gradients in the bulk of the NS implies that, in contrast with main-sequence stars, no internal convection is triggered. Therefore, the common self-sustained convective dynamos, which operate ubiquitously in different astrophysical scenarios (planets, stars, brown dwarfs…), are not operative in NS and cannot be responsible for the observed long-term presence of the strong magnetic fields.
Thus, the primary goal of NS cooling studies is to develop realistic evolutionary models that, when compared with thermal emission observations from NSs of varying ages, yield valuable insights into the chemical composition, magnetic field strength and configuration of the emitting regions, or the properties of denser matter deeper within the star (Page et al, 2004; Yakovlev and Pethick, 2004; Aguilera et al, 2008b, a; Yakovlev et al, 2008; Page, 2009; Tsuruta, 2009; Potekhin et al, 2015a, 2020).
@@
In contrast, the internal process of heat transfer from the NS core manifests as a thermal component originating from a significant portion of the star’s surface. For a few dozen isolated neutron stars, the X-ray spectra are dominated by this thermal component, reflecting the residual cooling after the NS is born, probably compensated by internal heating processes connected to magnetic field decay. In this case, when an estimate of the star’s age is available, one can examine the relationship between temperature and age, providing an indirect approach to investigate the physics of the NS interior through comparison of long-term cooling models with observations. Therefore, careful selection of sources is essential when testing these cooling models, focusing specifically on cases where internal processes govern the evolution. The sample of sources suitable for these studies is gradually expanding, primarily consisting of young ($t\lesssim 10^{4}$ yr) NSs (Viganò et al, 2013; Potekhin et al, 2015a, 2020; Marino et al, 2024), along with a small number of slightly older NSs ($\sim 10^{5}-10^{6}$ yr), detectable due to their proximity within a few hundred parsecs. These older sources are known as the magnificent seven, or X-ray dim isolated NSs (Haberl, 2007).
@@
NS cooling scenarios are divided into two types: 1) ”standard” or ”minimal” cooling, driven by modified Urca processes and possibly enhanced by core superconductivity or superfluidity, and 2) ”enhanced” cooling, with faster cooling at early ages, triggered by neutrino production due to direct Urca processes (Lattimer and Prakash, 2001), presence of hyperons (Anzuini et al, 2022), quark matter or meson condensates (Umeda et al, 1994). The Vela pulsar, with its unusually low temperature and thermal luminosity for its age, is a key example suggesting rapid neutrino emission linked to high central density or exotic matter. This concept, first proposed by Page and Applegate (1992), has awaited further data and confirmation for over three decades. Evidence of enhanced cooling existed for some isolated NSs, but uncertainties in spectral data, ages, distances, accretion history, and magnetic field dissipation in high-field stars ($B\gtrsim 10^{14}$ G) hindered definitive confirmation. Very recently, a new analysis of three young, nearby, extremely cold NSs, whose properties require enhanced cooling to align models with observations (Marino et al, 2024), allowing to establish some constraints on the EoS. Currently, there is compelling evidence for rapid cooling in at least some NSs, or possibly most of them, if one accounts for strong magnetic fields effects counteracting the fast cooling (see e.g. Section 7.2 in Aguilera et al, 2008a). In addition, similar conclusions about the need for fast cooling mechanisms have been reached by examining the cooling of NSs in Low Mass X-ray Binaries following an accretion phase (Mendes et al, 2022).
@@
In the same spirit, NS cooling has traditionally been used to establish constraints on the existence and properties of new particles beyond the Standard Model. By analyzing the thermal evolution of NSs, researchers can probe the parameter space of hypothetical particles that could enhance cooling via additional energy loss mechanisms. Specific examples include Hamaguchi et al (2018), who derived upper limits on the axion decay constant, Buschmann et al (2022) used the cooling behavior of young NSs to constrain the axion mass, and Gómez-Bañón et al (2024) tightened these constraints by incorporating structural effects in the envelope.
@@
To illustrate a particularly intriguing case, we briefly examine the debated observational claims regarding the rapid cooling of the NS linked to the Cassiopeia A (Cas A) supernova remnant. Initial detailed studies of its soft X-ray thermal spectrum, spanning multiple years, indicated a surface temperature drop of approximately 4% per decade (Heinke and Ho, 2010). This was interpreted as evidence of a superfluid transition in the NS core, leading to enhanced neutrino emission, which sparked significant interest in modeling and motivated further in-depth X-ray observations (Page et al, 2011; Shternin et al, 2011; Ho et al, 2015). However, the inferred cooling rate has gradually decreased to 2-3% as more data and improved instrumental calibrations became available (Posselt and Pavlov, 2018). The most recent reanalysis, using data up to 2020 and incorporating updated background and detector models, places the cooling rate between 1.6% and 2.2% per decade, depending on the treatment of the hydrogen column density (Wijngaarden et al, 2019). While this confirms that the NS in Cas A is cooling, the revised (slower) rate challenges some of the earlier interpretations that required very strong neutrino emissivity. More recently, a thorough and critical revision has raised some doubts about possible systematic biases or imperfect calibrations which further challenge the reliability of the previously inferred cooling rate values (Posselt and Pavlov, 2022). This highlights the intricate nature of both theoretical modeling and data analysis challenges, demonstrating the dynamic and evolving character of this research field.
@@ 2.1 Neutron star structure
Figure 1: Structure and composition of a $1.4\,M_{\odot}$ NS, with SLy EoS. The plot shows, as a function of density from the outer crust to the core, the following quantities: mass fraction in the form of nuclei $X_{h}$ (blue dot-dashed line), the fraction of electrons per baryon $Y_{e}$ (black dashes), the fraction of free neutrons per baryon $Y_{n}$ (red dashes), the atomic number $Z$ (dark green triple dot-dashed), the mass number AA (cyan long dashes), radius normalized to $R$ (pink solid), and the corresponding enclosed mass normalized to the star mass (green solid).
In Fig. 1 we show a typical profile of a NS, obtained with the EoS SLy4 (Douchin and Haensel, 2001), which is among the realistic EoS supporting a maximum mass compatible with the observations, $M_{\max}\sim 2.0$$2.2\,M_{\odot}$ (Demorest et al, 2010; Antoniadis et al, 2013; Margalit and Metzger, 2017; Ruiz et al, 2018; Radice et al, 2018; Cromartie et al, 2019). We show the enclosed radius and mass, and the fractions of the different components, as a function of density, from the outer crust to the core. For densities $\rho\gtrsim 4\times 10^{11}~\mbox{g~cm${}^{-3}$}$, neutrons drip out of the nuclei and, for low enough temperatures, they would become superfluid. Note that the core contains about 99% of the mass and comprises 70–90% of the star volume (depending on the total mass and EoS). Envelope and atmosphere are not represented here. For a more detailed discussion, see e.g., Haensel et al (2007); Potekhin et al (2015b).
@@ 2.2 Heat transfer equation
Spherical symmetry was also assumed in most NS cooling studies during the 1980s and 1990s. However, in the 21st century, the unprecedented amount of data collected by soft X-ray observatories such as Chandra and XMM-Newton supports that most nearby NSs whose thermal emission is visible in the X-ray band of the electromagnetic spectrum show some anisotropic temperature distribution (Haberl, 2007; Posselt et al, 2007; Kaplan et al, 2011). This observational evidence made clear the need to build multi-dimensional models and gave a new impulse to the development of the cooling theory including multidimensional effects (Geppert et al, 2004, 2006; Page et al, 2007; Aguilera et al, 2008b, a; Viganò et al, 2013; Beznogov et al, 2023). The cooling theory builds upon the heat transfer equation, which includes both flux transport and source/sink terms.
Figure 2: Contributions to the specific heat from neutrons (red dashes), protons (green dot-dashed), electrons (blue dots), and ions (black solid line) as a function of density, from the outer crust to the core, and for different temperatures in each panel (as indicated). The superfluid models employed here are the same as in (Ho et al, 2012). The plots refer to the representative NS shown in Fig. 1.
@@
$c_{\mathrm{v}}\,\frac{\partial(T\mathrm{e}^{\nu})}{\partial t}+\vec{\nabla}\cdot(\mathrm{e}^{2\nu}\vec{F})=\mathrm{e}^{2\nu}(H-Q)~,$ (4)
@@
$\vec{F}=-\mathrm{e}^{-\nu}\hat{\kappa}\cdot\vec{\nabla}(\mathrm{e}^{\nu}T)~,$ (5)
@@
$\frac{\kappa^{\parallel}}{\kappa^{\perp}}\approx 1+(\omega_{B}^{e}\tau_{e})^{2}~,$ (6)
@@
To understand the role of anisotropy in strong magnetic fields, we can examine electron conductivity while neglecting quantizing effects. The heat flux, as derived by (Pérez-Azorín et al, 2006), is expressed in the compact form:
$\vec{F}=-e^{-\nu}\kappa^{\perp}\left[\vec{\nabla}(e^{\nu}T)+(\omega_{B}^{e}\tau_{e})^{2}(\vec{b}\cdot\vec{\nabla}(e^{\nu}T))\vec{b}+\omega_{B}^{e}\tau_{e}(\vec{b}\times\vec{\nabla}(e^{\nu}T))\right]~,$ (7)

### v2

@@ Abstract
The strong magnetic fields of neutron stars are closely linked to their observed thermal, spectral, and timing properties, such as the distribution of spin periods and their derivatives. To understand the evolution of astrophysical observables over time, it is essential to develop robust theoretical frameworks and numerical models that solve the coupled thermal and magnetic field evolution equations, incorporating detailed microphysics such as thermal and electrical conductivities and neutrino emission rates. These efforts are key to uncovering how the strength and geometry of magnetic fields change with age, ultimately shedding light on the diverse phenomenology of neutron stars. In this review, we outline the fundamental theory underlying magneto-thermal evolution models, with an emphasis on numerical methods and a comprehensive set of benchmark tests intended to guide current and future code development. We revisit established results from axisymmetric simulations, highlight recent progress in fully three-dimensional models, and offer a perspective on the anticipated developments in this rapidly evolving field.
@@ 1 Introduction
Regarding observation, NSs were first discovered as rotation-powered radio pulsars (standing for pulsating stars, due to their periodic signal). These so-called standard pulsars constitute the most numerous class of known NSs, approaching four thousand identified members. [^1] See the online Australia Telescope National Facility pulsar catalog, http://www.atnf.csiro.au/research/pulsar/psrcat/ The number is continuously increasing thanks to the progress in wide-field low-frequency radio surveys and the use of the new generation of high-sensitivity radio interferometers, like LOFAR (van Haarlem, 2013) and the soon-available Square Kilometre Array Observatory (SKAO). The latter is foreseen to be able to detect many thousands of regular pulsars.
@@
To a lesser extent, NSs have also been observed in $X$-rays (about one hundred NSs so far), as persistent or transient sources, and/or as $\gamma$-ray pulsars (340 in the recent third Fermi-LAT pulsar catalog, Smith et al. 2023). The origin of this high-energy radiation is typically non-thermal, originated by particle acceleration (synchro-curvature emission, Zhang and Cheng 1997; Viganò et al. 2015) or Compton up-scattering of lower-energy photons by the particles composing the magnetospheric plasma (Lyutikov and Gavriil, 2006). An exception is the soft X-ray thermal emission from the surface, observed in only a few dozen, mostly young, neutron stars (Potekhin et al., 2015a); despite its rarity, it is highly relevant to this review.
@@
A particularly intriguing class of isolated NSs are the magnetars (Mereghetti et al., 2015; Turolla et al., 2015; Kaspi and Beloborodov, 2017; Rea and De Grandis, 2026), relatively slow rotators with typical spin periods of several seconds and ultra-strong timing-inferred magnetic fields ($10^{13}$$10^{15}$ G). In most cases, they show a relatively high persistent (i.e., constant over many years) X-ray luminosity ($L_{x}\approx 10^{33}$$10^{35}$ erg/s), well exceeding their rotational energy losses, in contrast with radio (standard) and $\gamma$-ray pulsars. This leads to the conclusion that the main source of energy is provided by the strong magnetic field, instead of rotational energy. Magnetars are also identified for their complex transient phenomenology in high energy $X$-rays and $\gamma$-rays, including short (tenths of a second) bursts, occasional energetic outbursts with months-long afterglows (Rea and Esposito, 2011; Coti Zelati et al., 2018) and, much more rarely (only three observed so far), giant flares (Hurley et al., 1999; Palmer et al., 2005). During giant flares, the energy release is as large as $10^{46}$ erg in less than a second. The source of energy of such transient, violent behavior is also generally agreed to be of magnetic origin, as originally proposed in Thompson and Duncan (1995, 1996).
@@
Although isolated NSs have been historically differentiated in sub-classes, mostly based on observational grounds (detectability in $X$ and/or radio, transient vs. persistent properties, and presence/absence of pulsations), there is arguably no sharp boundary between classes, and the distributions of their physical properties, such as the inferred magnetic field, partially overlap. Indeed, the evidence accumulated in the last few decades has shown that the presence of a strong dipolar field, which can be reliably estimated from timing properties (see Sect. 5), is not a sufficient condition to trigger observable magnetar-like events. In contrast, a growing number of NSs with relatively low inferred surface dipolar magnetic fields have been observed showing magnetar-like activity. These include some normal pulsars (Rea et al., 2010, 2012, 2013, 2014; Gavriil et al., 2008; Göğüs et al., 2016; Lower et al., 2021; Uzuner et al., 2023), often referred to as low-field magnetars, but also other sources belonging to the sub-class of central compact objects (CCOs), a handful of young NSs surrounded by a supernova remnant, detectable due to a persistent, mostly non-pulsating X-ray emission (De Luca, 2017). It is now evident that the complex, non-linear dynamics of the internal magnetic field, coupled with its interaction with the magnetospheric plasma, is crucial for understanding the observed phenomena.
@@ 2 Neutron star cooling
The evolution of the temperature in a NS was theoretically explored even before the first detections, in the 1960s (Tsuruta, 1964). Today, NS cooling is the most widely accepted terminology for the research area studying how internal and surface temperature evolve as NSs age and their observable effects. We refer the interested reader to the introduction in the review by Potekhin et al. (2015b) for a thorough historical overview of the foundations of the NS cooling theory.
@@
The high thermal conductivity of the core rapidly leads to a nearly isothermal [^3] We employ the term “isothermal” in the relativistic sense, including metric corrections, as we describe below. state within the first year. After a few years, the gradual long-term cooling of residual heat proceeds slowly, on a timescale of $10^{5}$$10^{6}$ years, with temperature gradients essentially limited to the crust and envelope. These temperature gradients are primarily radial, but, as discussed in Sect. 2.2, they can also develop in the angular directions in the presence of strong magnetic fields, leading to surface temperature variations that influence the observed X-ray spectra. Notably, the absence of strong temperature gradients in the bulk of the NS implies that, in contrast with main-sequence stars, no internal convection is triggered. Therefore, the common self-sustained convective dynamos, which operate ubiquitously in different astrophysical scenarios (planets, stars, brown dwarfs…), are not operative in NS and cannot be responsible for the observed long-term presence of the strong magnetic fields.
Thus, the primary goal of NS cooling studies is to develop realistic evolutionary models that, when compared with thermal emission observations from NSs of varying ages, yield valuable insights into the chemical composition, magnetic field strength and configuration of the emitting regions, or the properties of denser matter deeper within the star (Page et al., 2004; Yakovlev and Pethick, 2004; Aguilera et al., 2008b, a; Yakovlev et al., 2008; Page, 2009; Tsuruta, 2009; Potekhin et al., 2015a, 2020).
@@
In contrast, the internal process of heat transfer from the NS core manifests as a thermal component originating from a significant portion of the star’s surface. For a few dozen isolated neutron stars, the X-ray spectra are dominated by this thermal component, reflecting the residual cooling after the NS is born, probably compensated by internal heating processes connected to magnetic field decay. In this case, when an estimate of the star’s age is available, one can examine the relationship between temperature and age, providing an indirect approach to investigate the physics of the NS interior through comparison of long-term cooling models with observations. Therefore, careful selection of sources is essential when testing these cooling models, focusing specifically on cases where internal processes govern the evolution. The sample of sources suitable for these studies is gradually expanding, primarily consisting of young ($t\lesssim 10^{4}$ yr) NSs (Viganò et al., 2013; Potekhin et al., 2015a, 2020; Marino et al., 2024), along with a small number of slightly older NSs ($\sim 10^{5}$$10^{6}$ yr), detectable due to their proximity within a few hundred parsecs. These older sources are known as the magnificent seven, or X-ray dim isolated NSs (Haberl, 2007).
@@
NS cooling scenarios are divided into two types: 1) “standard” or “minimal” cooling, driven by modified Urca processes and possibly enhanced by core superconductivity or superfluidity, and 2) “enhanced” cooling, with faster cooling at early ages, triggered by neutrino production due to direct Urca processes (Lattimer and Prakash, 2001), presence of hyperons (Anzuini et al., 2022), quark matter or meson condensates (Umeda et al., 1994). The Vela pulsar, with its unusually low temperature and thermal luminosity for its age, is a key example suggesting rapid neutrino emission linked to high central density or exotic matter. This concept, first proposed by Page and Applegate (1992), has awaited further data and confirmation for over three decades. Evidence of enhanced cooling existed for some isolated NSs, but uncertainties in spectral data, ages, distances, accretion history, and magnetic field dissipation in high-field stars ($B\gtrsim 10^{14}$ G) hindered definitive confirmation. Very recently, a new analysis of three young, nearby, extremely cold NSs, whose properties require enhanced cooling to align models with observations (Marino et al., 2024), allowing to establish some constraints on the EoS. Currently, there is compelling evidence for rapid cooling in at least some NSs, or possibly most of them, if one accounts for strong magnetic fields effects counteracting the fast cooling (see e.g. Sect. 7.2 in Aguilera et al., 2008a). In addition, similar conclusions about the need for fast cooling mechanisms have been reached by examining the cooling of NSs in Low Mass X-ray Binaries following an accretion phase (Mendes et al., 2022).
@@
In the same spirit, NS cooling has traditionally been used to establish constraints on the existence and properties of new particles beyond the Standard Model. By analyzing the thermal evolution of NSs, researchers can probe the parameter space of hypothetical particles that could enhance cooling via additional energy loss mechanisms. Specific examples include Hamaguchi et al. (2018), who derived upper limits on the axion decay constant, while Buschmann et al. (2022) used the cooling behavior of young NSs to constrain the axion mass, and Gómez-Bañón et al. (2024) tightened these constraints by incorporating structural effects in the envelope.
@@
To illustrate a particularly intriguing case, we briefly examine the debated observational claims regarding the rapid cooling of the NS linked to the Cassiopeia A (Cas A) supernova remnant. Initial detailed studies of its soft X-ray thermal spectrum, spanning multiple years, indicated a surface temperature drop of approximately 4% per decade (Heinke and Ho, 2010). This was interpreted as evidence of a superfluid transition in the NS core, leading to enhanced neutrino emission, which sparked significant interest in modeling and motivated further in-depth X-ray observations (Page et al., 2011; Shternin et al., 2011; Ho et al., 2015a). However, the inferred cooling rate has gradually decreased to 2–3% as more data and improved instrumental calibrations became available (Posselt and Pavlov, 2018). The most recent reanalysis, using data up to 2020 and incorporating updated background and detector models, places the cooling rate between 1.6% and 2.2% per decade, depending on the treatment of the hydrogen column density (Wijngaarden et al., 2019). While this confirms that the NS in Cas A is cooling, the revised (slower) rate challenges some of the earlier interpretations that required very strong neutrino emissivity. More recently, a thorough and critical revision has raised some doubts about possible systematic biases or imperfect calibrations which further challenge the reliability of the previously inferred cooling rate values (Posselt and Pavlov, 2022). This highlights the intricate nature of both theoretical modeling and data analysis challenges, demonstrating the dynamic and evolving character of this research field.
@@ 2.1 Neutron star structure
Figure 1: Structure and composition of a $1.4\,M_{\odot}$ NS, with SLy EoS. The plot shows, as a function of density from the outer crust to the core, the following quantities: mass fraction in the form of nuclei $X_{h}$ (blue dot-dashed line), the fraction of electrons per baryon $Y_{e}$ (black dashes), the fraction of free neutrons per baryon $Y_{n}$ (red dashes), the atomic number $Z$ (dark green triple dot-dashed), the mass number AA (cyan long dashes), radius normalized to $R$ (pink solid), and the corresponding enclosed mass normalized to the star mass (green solid). Note that the total neutron fraction (not shown here) varies continuously across the NS interior, unlike the fraction of free neutrons per baryon, $Y_{n}$.
In Fig. 1 we show a typical profile of a NS, obtained with the EoS SLy4 (Douchin and Haensel, 2001), which is among the realistic EoS supporting a maximum mass compatible with the observations, $M_{\max}\sim 2.0$$2.2\,M_{\odot}$ (Demorest et al., 2010; Antoniadis et al., 2013; Margalit and Metzger, 2017; Ruiz et al., 2018; Radice et al., 2018; Cromartie et al., 2020). We show the enclosed radius and mass, and the fractions of the different components, as a function of density, from the outer crust to the core. For densities $\rho\gtrsim 4\times 10^{11}~\mbox{g~cm${}^{-3}$}$, neutrons drip out of the nuclei and, for low enough temperatures, they would become superfluid. Note that the core contains about 99% of the mass and comprises 70–90% of the star volume (depending on the total mass and EoS). Envelope and atmosphere are not represented here. For a more detailed discussion, see e.g., Haensel et al. (2007); Potekhin et al. (2015b).
@@ 2.2 Heat transfer equation
Spherical symmetry was also assumed in most NS cooling studies during the 1980s and 1990s. However, in the 21st century, the unprecedented amount of data collected by soft X-ray observatories such as Chandra and XMM-Newton supports that most nearby NSs whose thermal emission is visible in the X-ray band of the electromagnetic spectrum show some anisotropic temperature distribution (Haberl, 2007; Posselt et al., 2007; Kaplan et al., 2011). This observational evidence made clear the need to build multi-dimensional models and gave a new impulse to the development of the cooling theory including multidimensional effects (Geppert et al., 2004, 2006; Page et al., 2007; Aguilera et al., 2008b, a; Viganò et al., 2013; Beznogov et al., 2023). The cooling theory builds upon the heat transfer equation, which includes both flux transport and source/sink terms.
Figure 2: Contributions to the specific heat from neutrons (red dashes), protons (green dot-dashed), electrons (blue dots), and ions (black solid line) as a function of density, from the outer crust to the core, and for different temperatures in each panel (as indicated). The superfluid models employed here are the same as in Ho et al. (2012). The plots refer to the representative NS shown in Fig. 1.
@@
$c_{\mathrm{v}}\,\frac{\partial(T\mathrm{e}^{\nu})}{\partial t}+\vec{\nabla}\cdot(\mathrm{e}^{2\nu}\vec{F})=\mathrm{e}^{2\nu}(H-Q)\,$, (4)
@@
$\vec{F}=-\mathrm{e}^{-\nu}\hat{\kappa}\cdot\vec{\nabla}(\mathrm{e}^{\nu}T)\,$, (5)
@@
$\frac{\kappa^{\parallel}}{\kappa^{\perp}}\approx 1+(\omega_{B}^{e}\tau_{e})^{2}\,$, (6)
@@
To understand the role of anisotropy in strong magnetic fields, we can examine electron conductivity while neglecting quantizing effects. The heat flux, as derived by Pérez-Azorín et al. (2006), is expressed in the compact form:
$\vec{F}=-e^{-\nu}\kappa^{\perp}\left[\vec{\nabla}(e^{\nu}T)+(\omega_{B}^{e}\tau_{e})^{2}(\vec{b}\cdot\vec{\nabla}(e^{\nu}T))\vec{b}+\omega_{B}^{e}\tau_{e}(\vec{b}\times\vec{\nabla}(e^{\nu}T))\right]\,$, (7)